OpenStudy (anonymous):

2 + (a + b) = (2 + a) + b is an example of which algebraic property? A. Associative Property of Addition B. Commutative Property of Addition C. Symmetric Property D. Distributive Property

OpenStudy (kinggeorge):

Do you have a guess?

OpenStudy (anonymous):

Common, this is LOL question :P

OpenStudy (anonymous):

*comm'on

OpenStudy (anonymous):

I think it's "Distributive Property" but I'm not sure...

OpenStudy (anonymous):

HInt: the answer is the first letter of the object which fell on Newton's head :P

OpenStudy (anonymous):

Hmm... Is it 'C'? No. Just kidding. Okay, thanks.

OpenStudy (anonymous):

It was a Capple that fell on his head, right? :D

OpenStudy (anonymous):

LOLOL, that apple changed the size of our physics textbooks :P

OpenStudy (accessdenied):

Associating basically means putting things into a group. Associative property deals with grouping things together. (a + b) + c = a + (b + c) Commuting basically means to change something or travel. Commutatie property deals with changing the order of the expressions being added/multiplied. a + b = b + a Symmetry is the property that it looks the same folded across some line. The symmetric property of equality is to flip the equality around so that the first side is on the other one. a = b -> b = a Distributing means to send things to a large group. Distributive property is when you have a number times a sum, you multiply each term in the sum by that number. a(b + c) = ab + ac "a distributed to b and c" - maybe that'd help distinguish which is which, since their names basically tell the story themselves.

OpenStudy (anonymous):

@AccessDenied , good info for beginners :)